# Conductivity of concentrated salt solutions

### v4

@@ Abstract
The conductivity of concentrated salt solutions has posed a real puzzle for theories of electrolytes. Despite a quantitative understanding of dilute solutions, an analytical theory for concentrated ones remains a challenge almost a century, although a number of parameters and effects incorporated into theories increases with time. Here we show that the conductivity of univalent salt solutions can be perfectly interpreted using a simple model that relies on a modified mean-field description of electrostatic effects and on a classical approach to calculating colloid electrophoresis. We derive a compact equation, which predicts that the ratio of conductivity at a finite concentration to that at an infinite dilution is the same for all salt, if it is plotted against a product of the harmonic mean of ion hydrodynamic radii and the square root of concentration. Our equation fits very well the data for inorganic salts (up to a few mol/l), although at a very high dilution the relaxation correction seems necessary.
@@ I Introduction
The physical origin of ionic conductivity is more or less clear. If an electric field $E$ is applied, ions migrate relative to a solvent by generating an electric current of density $J=KE$ (Ohm’s law). The migration speeds of ions are given by
@@
The quantitative understanding of ion mobilities is a challenging problem that has been addressed over nearly a century and by many groups, which is often termed a central issue of chemical physics [6, 7]. The simplest expression can be derived by postulating the Stokes resistance to the ion propulsion
@@
where $\eta$ is the dynamic viscosity of a solvent and $R_{h}=2(R_{+}R_{-})/(R_{+}+R_{-})$ is the harmonic mean of hydrodynamic radii. However, experiments on a conductivity show that $M$ and $K$ are generally smaller than predicted by (4) and this discrepancy augments on increasing salt concentration [8, 1]. Chemists have long used this fact to infer a degree of dissociation (or an ion pairing). Physicists might view this simply as unreliability of hydrodynamic arguments based on the Stokes force.
@@
There is a large literature describing attempts to provide a satisfactory theory of electrolyte conductivity. One of the first systematic treatments of the influence of $n_{\infty}$ on conductivity was contained in a remarkable paper by Onsager [9] who clarified a mechanism for the conductivity reduction, which involved ionic electrophoresis and relaxation, and showed that at a very high dilution the correction to $M_{0}$ depends on the square root of the concentration. In efforts to better understand the connection between the ion mobilities and salt concentration many authors extended the theory, but failed to come to grips with the conductivity at high salt [10]. However, most electrolyte solutions in nature and various applications are concentrated. Thus, salt concentration in human blood plasma is about 0.15 mol/l, in the Atlantic Ocean it is $\simeq 0.6$ mol/l, Li-ion batteries and supercapacitor electrolytes are usually of concentration $1-2$ mol/l, reference electrolytes of pH-meters and glass micropipette electrodes - of concentration 3 mol/l.
@@
Recent attempts at improvements on the theory has been focused on a more sophisticated description of electrostatic and hydrodynamic effects. Various techniques, such as the mean spherical approximation [11, 6], mode-coupling theory [12, 13], density functional approach [12, 14], and more [15, 16, 17, 7, 18] have been employed. These publications involve additional parameters and contributions, and mostly rely on numerical calculations. This makes them difficult to use and limits prediction capabilities, but simple analytical formulas applied for concentrated solutions have not yet been obtained, although there have been some attempts to deduce them from microscopic theories. Useful analytical solution is known [19], but expressed in terms of special functions and can hardly be called simple. Avni et al. [14] derived an approximate conductivity expression, but do not recognize its inapplicability at molar concentrations due to non physical behavior of electrophoretic and relaxation terms [20]. Recently we suggested a completely different line to attack the ionic electrophoresis problem [21]. Assuming cations and anions of equal size, we have carried calculations of an electrophoretic term and derived an equation for a corresponding reduction in relative conductivity, $K/K_{0}$, that applies even at a very low dilution. However, this paper has not given sufficient attention to the relaxation issue, which is traditionally invoked in the conductivity theories. It is also, in practice, unlikely that both cations and anions have exactly the same size, so the status of our theory for real salt solutions remains somewhat obscure.
@@
The present paper extends and generalizes the prior analysis [21] to solutions composed of inorganic ions of unequal radius $R_{+}\neq R_{-}$. We derive a simple equation for an electrophoretic contribution to conductivity valid up to a few mol/l and argue that relative conductivities, $K/K_{0}$, of all inorganic salts plotted against $R_{h}\sqrt{n_{\infty}}$ would collapse into a single curve. This conclusion is supported by providing a comparison with data for several standard salts, which shows that our theory is in excellent agreement with experiment up to concentrations of a few molars. We also demonstrate that although the relaxation effect affects a small decrement to a relative conductivity at high dilution, it becomes negligibly small in concentrated solutions, where the reduction in $K/K_{0}$ is significant.
@@ II General considerations
We consider a bulk 1:1 salt solution assuming that the description of its global static properties can be restricted to a mean-field Poisson-Boltzmann theory. The dimensionless electrostatic potential around each ion $\phi(z)=e\Phi(z)/(k_{B}T)$, where $k_{B}$ is the Boltzmann constant and $T$ is a temperature, represents a continuous function that depends on all other ions. The Debye screening length of a solution, $\lambda_{D}=\left(8\pi\ell_{B}n_{\infty}\right)^{-1/2}$, is defined as usually with the Bjerrum length, $\ell_{B}=\dfrac{e^{2}}{\varepsilon k_{B}T}$, where $\varepsilon$ is the solvent permittivity. By analyzing the experimental data it is more convenient to use the concentration $c_{\infty}[\rm{mol/l}]$, which is related to $n_{\infty}[\rm{m^{-3}}]$ as $n_{\infty}\simeq N_{A}\times 10^{3}\times c_{\infty}$, where $N_{A}$ is Avogadro’s number. The Bjerrum length of water at $T\simeq 298$ K is equal to about $0.7$ nm leading to a useful formula for 1:1 electrolyte
@@
Thus upon increasing $c_{\infty}$ from $10^{-5}$ to $3$ mol/l the screening length is reduced from about 100 down to 0.18 nm.
@@
We recall that the mean-field theory neglects correlations and finite sizes of ions. By contrast, to describe the dynamic response to an external field, the ion hydrodynamic radii $R_{\pm}$ should be taken into account. Inorganic ions have hydrodynamic radii from 0.1 to 0.3 nm [22], and we present their values for some univalent electrolytes in Table 1, together with harmonic means $R_{h}=2R_{+}R_{-}/(R_{+}+R_{-})$ and $N=R_{+}/R_{-}$. These apparent contradictions indicate that in reality, ions are neither point-like nor particles of finite size, exhibiting the properties of the former or the latter only depending on the circumstances. This can be referred to as the point-particle duality of ions (by analogy to the familiar concept of wave-particle duality of electrons in quantum mechanics). Such a duality can be introduced, if an electrolyte solution is modelled as a colloidal suspension of so-called central ions representing isolating particles of radius $R_{\pm}$ with a constant surface charge density ±$\pm e/(4\pi R_{\pm}^{2})$ [21]. Such a sphere induces the same outer field as a point charge ±$\pm e$ located at its center would, and the salt-dependent surface potential $\phi_{s\pm}$ is established self-consistently. By contrast, the electrostatic diffuse layers (EDLs or clouds) are treated on the classical mean-field level and formed by point-like ionic species, thus their extension is defined by $\lambda_{D}$. Two further comments should be made: (i) Since the model prevents point-like EDL ions from penetrating inside the sphere, the need to invoke an additional electrostatic cut-off [14] is removed. (ii) The approach neglects electrostatic correlations inside the cloud, but provides positional correlations between central ions, although on average they are randomly distributed.
@@
Since the electrostatics of ions involves two length scales, it is convenient to introduce their dimensionless radii $\varrho_{\pm}=R_{\pm}/\lambda_{D}$. Note that in the range of $c_{\infty}$ below 3 mol/l we consider here, the values of ϱ±subscriptitalic-$\varrho_{\pm}$ remain smaller than unity or very close to it. Say, if we set $c_{\infty}=2$ mol/l, then for a largest ion (Li+) in Table 1 we obtain $\varrho_{+}\simeq 1$, but for a smallest (I-) we get $\varrho_{-}\simeq 0.5$.
@@
When an electric field $E$ is applied, an electro-osmotic flow around central ions, referred below simply to as ions, is induced. The electroosmosis takes its origin in the EDL, where a tangential electric field generates a force that sets the fluid in motion. The emergence of this flow in turn provides hydrodynamic stresses that cause the propulsion of ions with a velocity given by Eq.(1) as sketched in Fig. 1. The space between clouds of two ions is electro-neutral, and this region is force-free (no flow) provided they do not overlap. Thus, the problem of ionic electrophoresis can be treated on a single-particle level.
@@
We return to the importance of the relaxation later, by focusing now on electrophoresis solely. This contribution to a conductivity is traditionally calculated assuming that the surrounding cloud remains undisturbed during the migration. The electrophoretic mobilities of these ions can be defined using the known expression for an electro-osmotic mobility [23, 24] taken with the opposite sign
@@
where $Z_{\pm}$ are the electrokinetic or zeta potentials of cations and anions, and $\zeta_{\pm}=eZ_{\pm}/k_{B}T$. In such a definition the geometry factor is hidden inside $\zeta$. Note that by introducing the dimensional velocity $v_{\pm}=\dfrac{4\pi\eta\ell_{B}}{eE}V_{\pm}$, one can rewrite (1) as $v_{\pm}=\zeta_{\pm},$ which points clearly that the electrophoretic speed of ions is set by their zeta potentials.
III Results and discussion
Zeta potentials of small inorganic ions are given by [21]
@@ III Results and discussion
where $\mathcal{F_{\pm}}$ are the special functions derived by Henry [25]
@@
±$\mathcal{F_{\pm}}\simeq\dfrac{2}{3}\left[1+\left(\dfrac{\varrho_{\pm}}{4}\right)^{2}\right]$ (9)
[…]
$\zeta_{\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}(1+\varrho_{\pm})}.$ (10)
Clearly, the upper bound for the zeta potential is attained when $\rho_{\pm}\to 0$ (an infinite dilution)
@@
Figure 2: Zeta potentials $\zeta_{\pm}$ computed for K+, Li+, Cl-, I- (solid curves from top to bottom). Dashed lines are obtained using Eq. (11). Symbols show calculations from Eq. (10).
Figure 2 includes computed zeta potentials of several ions. For these examples we take from Table 1 cations (K+, Li+) and anions (I-, Cl-) of smallest and largest radius. Calculations are made using the first equality in (7) with $\phi_{s\pm}$ found numerically from the nonlinear Poisson-Boltzmann equation and $\mathcal{F_{\pm}}$ calculated from Eq. (8). The straight lines corresponding to $\zeta_{0\pm}$ given by (11) describe perfectly the distinct plateau regions at very low concentrations. On increasing concentration, however, the absolute values of $\zeta_{\pm}$ decrease. Also included are calculations from (10). It can be seen that their fit to numerical data is extremely good in the whole concentration range.
Refer to captionRefer to captionRefer to caption
Figure 3: $K_{\infty}/K_{\infty}^{0}$ as a function of $c_{\infty}$ for KBr, NaCl and LiI aqueous solution calculated from Eqs. (13), (14) and (15) (solid, dash-dotted and dashed curves). Open and filled circles indicate experimental data from [27] and [28], open and filled squares show data from [29] and [30].
@@
where $M_{0}$ and $K_{0}$ are given by Eqs. (4). Figure 3 shows $K/K_{0}$ calculated from Eq. (13) as a function of $c_{\infty}$. The calculations are compared with experimental data for three standard salts [27]. As seen from Table 1, for these examples $N$ varies from 1.1 for KBr, but can be larger than 2 (e.g. for LiI). An overall conclusion from this linear scale plot is that the theoretical curve for $K/K_{0}$ is well consistent with experiment. At large concentrations Eq.(13) predicts either slightly lower (KBr), but still fits data quite well. When $N=1$, we get $\varrho_{h}=R_{h}/\lambda_{D}=2\varrho_{+}/(1+N)=\varrho_{+}$, and Eq. (13) reduces to
@@
Eq. (14) predicts a monotonic decrease in relative (electrophoretic) conductivity with $\varrho$ as it should be, and $K/K_{0}\to 0$ as $\varrho\to\infty$. For $\varrho_{h}\ll 1$ it reduces to the famous Onsager formula for the electrophoretic effect

### v5

@@ Abstract
The conductivity of concentrated salt solutions has posed a real puzzle for theories of electrolytes. Despite a quantitative understanding of dilute solutions, an analytical theory for concentrated ones remains a challenge for almost a century, although a number of parameters and effects incorporated into theories increases with time. Here we show that the conductivity of univalent salt solutions can be perfectly interpreted using a simplest model that relies on a modified mean-field description of electrostatic interactions and on a classical approach to calculating colloid electrophoresis. We derive a compact equation, which predicts that the ratio of conductivity to that at an infinite dilution is the same for all salt and depends only on a product of the harmonic mean of ion hydrodynamic radii and the square root of concentration. Our equation fits very well the data for inorganic salts (up to a few mol/l), although at a very high dilution the relaxation correction seems necessary.
@@ I Introduction
The physical origin of ionic conductivity is more or less clear. If a direct electric field $E$ is applied, ions migrate relative to a solvent with a constant speed towards one of the electrodes. Such a migration results in an electric current, which density obeys Ohm’s law, $J=KE$, provided the field is weak enough. The speeds of ions are given by
@@
The quantitative understanding of ion mobilities is a challenging problem that has been addressed over nearly a century and by many groups, which is often termed a central issue of chemical physics [6, 7]. The simplest expression for univalent ions can be derived by postulating the Stokes resistance to the ion propulsion
@@
where $\eta$ is the dynamic viscosity of a solvent and $R_{h}=2(R_{+}R_{-})/(R_{+}+R_{-})$ is the harmonic mean of hydrodynamic radii. However, experiments on a conductivity show that $M$ and $K$ are generally smaller than predicted by (4) and this discrepancy augments on increasing salt concentration [8, 1]. Chemists have long (since Arrhenius) used this fact to infer a degree of dissociation (or an ion pairing). Physicists might view this simply as unreliability of hydrodynamic arguments based on the Stokes drag force.
@@
There is a large literature describing attempts to provide a satisfactory theory of electrolyte conductivity. One of the first systematic treatments of the influence of $n_{\infty}$ on conductivity was contained in a remarkable paper by Onsager [9] who argued that the relative conductivity reduction $K/K_{0}$ is attributed to two contributions. One (and the main) is ionic electrophoresis, which is the migration of ion surrounded by equilibrium screening cloud. Second is a so-called relaxation correction that depends on the cloud distortion produced by the motion in the external field. Onsager showed that at a very high dilution $K/K_{0}$ decreases linearly with the square root of the concentration. In efforts to better understand the connection between the ion mobilities and salt concentration many authors extended the theory, but failed to come to grips with the conductivity at high salt [10]. However, most electrolyte solutions in nature and various applications are concentrated. Thus, salt concentration in human blood plasma is about 0.15 mol/l, in the Atlantic Ocean it is 0.6$\simeq$ 0.6≃ 0.6 mol/l, Li-ion batteries and supercapacitor electrolytes are usually of concentration 121-21 - 2 mol/l, reference electrolytes of pH-meters and glass micropipette electrodes - of concentration 3 mol/l.
@@
Recent attempts at improvements on the theory have been focused on a sophisticated description of so-called electrostatic and hydrodynamic effects based on microscopic concepts. Various techniques, such as the mean spherical approximation [11, 6], mode-coupling theory [12, 13], density functional approach [12, 14], and more [15, 16, 17, 7, 18] have been employed. These publications involve additional parameters and contributions, and mostly rely on numerical calculations. This makes them difficult to use and limits prediction capabilities, but simple analytical formulas that apply for concentrated solutions appear to be missing, although there have been some attempts to deduce them from microscopic theories. A useful analytical solution is known [19], but expressed in terms of special functions and can hardly be called simple or compact. A more recent approximate expression [14] is inapplicable at molar concentrations due to non physical behavior of both electrophoretic and relaxation terms [20]. Thus, in spite of importance for many fields, a theory that has the merit of yielding useful (approximate) analytical result for the conductivity of concentrated salt solutions as well as makes strong appeal to macroscopic arguments is still absent.
We also gained the impression that researchers in one branch appear to be unaware of closely related work in other branches of physics and chemistry. Indeed, most conductivity publications have either ignored or failed to make direct connection with the theory of electrophoresis developed in colloid science and fluid mechanics. Thus, already in 1921 Smoluchowski [21] made clear that the electrophoretic mobility is proportional to the electrostatic surface potential $\Phi_{s}$, and not, as one might imagine upon first thought, to the charge of the particle. Since in all existing theories ions are idealized as hard spheres of radius $a_{\pm}$, this conclusion holds for ions too, at least within the framework of such a colloid representation. However, the significance of the Smoluchowski paper does not seem to be recognized in the modern conductivity literature. Recently we have drawn attention to this drawback of the conductivity theories and have carried calculations of $\Phi_{s}$ and then of the so-called zeta potential of ions $Z\propto\Phi_{s}$, which defines their electrokinetic mobility [22]. Assuming cations and anions of equal hydrodynamic radius, $R_{+}=R_{-}$, and using $a=R_{\pm}$, we have derived an equation for $K/K_{0}$, that applies even at a very low dilution and is more accurate and compact than prior approximate expression [14]. However, our paper has not given attention to the relaxation issue, which is traditionally invoked in the conductivity theories. It is also, in practice, unlikely that both cations and anions have exactly the same size, so the status of the theory for real salt solutions still remains somewhat obscure.
@@
The present paper concerns with solutions composed of inorganic ions of unequal radius $R_{+}\neq R_{-}$. We revisit the issue of calculation of ion zeta potentials and propose an even much simpler expression than derived before [22], which appears to be more accurate and applies at a larger concentration range. We then use it to derive the simplest conductivity equation valid up to several mol/l, and argue that relative conductivities, $K/K_{0}$, of all inorganic salts plotted against $R_{h}\sqrt{n_{\infty}}$ would collapse into a single curve. This conclusion is supported by providing a comparison with data for several standard salts, which shows that our theory is in excellent agreement with experiment up to concentrations of a few molars. We also demonstrate that although the relaxation effect affects a small decrement to a relative conductivity at high dilution, it becomes negligibly small in concentrated solutions, where the reduction in $K/K_{0}$ is significant.
@@ II General considerations
Figure 1: Sketch of the charge distribution in the (a) real electrolyte solution, (b) Debye-Hückel classical model of finite size ions with the charges located at their center, and (c) model of ions with neutral interior, but constant surface charge densities.
We consider a bulk 1:1 salt solution of concentration $n_{\infty}[\rm{m^{-3}}]$ assuming that the description of its global static properties can be restricted to a mean-field theory. Our approach is generally based on the fundamental ideas of Debye and Hückel [23]. Namely, an electrolyte solution with randomly distributed ions [shown in Fig. 1(a)] is idealized as a colloidal suspension of so-called central (positive and negative) ions representing nonconducting hard spheres of finite radii $a_{\pm}$ carrying unit charges ±$\pm e$ at their center. Compensating charges of the opposite sign and equal magnitude stay in the neighborhood of the central ions. These are the so-called ionic clouds or electrostatic diffuse layers (EDLs). EDLs are treated as composed by point-like ionic species, and their extension is of the order of the Debye screening length of a solution, $\lambda_{D}=\left(8\pi\ell_{B}n_{\infty}\right)^{-1/2}$, where $\ell_{B}=\dfrac{e^{2}}{\varepsilon k_{B}T}$ the Bjerrum length of the solvent, $\varepsilon$ is the solvent permittivity, $k_{B}$ is the Boltzmann constant, and $T$ is a temperature. This classical model is illustrated in Fig. 1(b).
Note that by analyzing the experimental data it is more convenient to use the concentration $c_{\infty}[\rm{mol/l}]$, which is related to $n_{\infty}$ as $n_{\infty}\simeq N_{A}\times 10^{3}\times c_{\infty}$, where $N_{A}$ is Avogadro’s number. The Bjerrum length of water at $T\simeq 298$ K is equal to about 0.70.70.7 nm leading to a useful formula for 1:1 electrolyte
@@
Thus upon increasing $c_{\infty}$ from $10^{-5}$ to 444 mol/l the screening length (and the cloud extension) is reduced from about 100 down to 0.15 nm.
Some further comments should be made. The fundamental insight of the Debye and Hückel approach is to realize that although the average potential inside the electrolyte solution is zero, there are strong positional correlations between the oppositely charged central ions [24]. Indeed, if some cation is selected as central, then the position of a central anion cannot be arbitrary anymore, and their local ordering inevitably occurs, although on average they are randomly distributed. Another important insight is to realize what we would term the point-particle duality of ions (by analogy to the familiar concept of wave-particle duality of electrons in quantum mechanics): ions can be either point-like or exhibit properties of the particles of finite size depending on the circumstances.
This classical model that is widely used to interpret equilibrium properties of electrolyte solutions will provide the starting point for our discussion of migration of central ions under an applied electric field.
In the interior of the ions of radii $a_{\pm}$ there are no other charges except for the fixed at the centers, and the inner electrostatic potentials satisfy the Laplace equation, $\nabla^{2}\phi_{\pm}=0$, where $\phi_{\pm}=e\Phi_{\pm}/(k_{B}T)$ are the dimensionless electrostatic potentials. This inner field does not impact the electrophoresis, which depends only on the surface potentials $\phi_{s\pm}$ with respect to the solution far from them (with the zero average potential), as stressed in Sec. I. The distribution of electrostatic potentials around the central ions can be obtained by application of the non-linear Poisson-Boltzmann equation [25]:
$\nabla^{2}\phi_{\pm}=\dfrac{1}{r^{2}}\dfrac{d}{dr}\left(r^{2}\dfrac{d\phi_{\pm}}{dr}\right)=\frac{\sinh\phi_{\pm}}{\lambda_{D}^{2}},$ (6)
where $r$ is the distance from the centers of the corresponding spheres. The outer potentials would be the same as for the spheres with a constant surface charge density ±$\pm e/(4\pi a_{\pm}^{2})$. and decay from $\phi_{s\pm}$ to zero as $r\to\infty$, which is the first boundary condition that should be imposed to integrate (6). While this condition is applied at infinity, as we already discussed above, in reality the EDLs extend only to distances of the order of $\lambda_{D}$, and thus potentials in fact vanish beyond the clouds. The surface potentials are established self-consistently and salt-dependent. It is therefore admissible for electrophoretic calculations to replace the standard classical model shown in Fig. 1(b) by spheres with a constant densities of a surface charge and neutral interior [see Fig. 1(c)].
@@
What is still missing and unknown are the radii $a_{\pm}$, which are the cut-off distances from the selected point charges, where there are no other ions. Some approaches to quantify them have been proposed [14, 6], but nevertheless, not developed enough to describe accurately the conductivity data. It is, however, well understood that to describe the dynamic response to an external field, the ion hydrodynamic radii $R_{\pm}$ should be taken into account. Inorganic ions have hydrodynamic radii from 0.1 to 0.3 nm [26], and we present their values for some univalent electrolytes in Table 1, together with harmonic means $R_{h}=2R_{+}R_{-}/(R_{+}+R_{-})$ and $N=R_{+}/R_{-}$. Since the introduction of $R_{\pm}$ inevitably prevents point-like EDL ions from penetrating inside the sphere, the need to invoke an additional electrostatic cut-off [14] is removed. So, below we simply postulate $a_{\pm}=R_{\pm}$ and impose the electrostatic boundary condition $\phi^{\prime}_{\pm}(r=R_{\pm})=\mp\ell_{B}/R_{\pm}^{2}$ to integrate Eq. (6) at this surface.
@@
Since the electrostatics of ions involves two length scales, it is convenient to introduce their dimensionless radii $\varrho_{\pm}=R_{\pm}/\lambda_{D}$. Note that in the range of $c_{\infty}$ below 2 mol/l the values of $\varrho_{\pm}$ remain smaller than unity or very close to it. Say, if we set $c_{\infty}=2$ mol/l, then for the largest ion (Li+) in Table 1 we obtain $\varrho_{+}\simeq 1.1$, but for the smallest (I-) we get $\varrho_{-}\simeq 0.5$. However, when $c_{\infty}=4$ mol/l the same ions give $\varrho_{+}\simeq 1.6$ and $\varrho_{-}\simeq 0.7$.
@@
When an electric field $E$ is applied, an electro-osmotic flow around central anions and cations, referred below simply to as ions, is induced. The electroosmosis takes its origin in their EDLs, where a tangential electric field generates a force that sets the fluid in motion. The emergence of this flow in turn provides hydrodynamic stresses that cause the propulsion of ions with a velocity given by Eq. (1) as sketched in Fig. 2. If the clouds of two ions do not overlap, the space between EDLs remains electro-neutral, and, therefore, this region is force-free. Consequently, in this outer region there is no fluid flow. Thus, within our model the problem of ionic electrophoresis can be treated on a single-particle level provided the concentration is not too high (below 4 mol/l or so). This, of course, significantly simplifies the problem. However, it would be worthwhile to stress that even the problem for a single colloidal sphere, the so-called “standard electrokinetic model” [27], can in general be solved only numerically [since analytical solution of Eq. (6) has not been found yet].
In the standard considerations of the electrophoretic migration, an essential supposition is normally that the radial electric field due to the double layer and the externally applied field may be simply superimposed on each other [28]. The Péclet number is assumed to be small, so that the convective fluxes of ions are ignored. If so, the system of electrokinetic equations is decoupled and the potential obeys the Poisson-Boltzmann equation [29], implying that the EDLs are undisturbed by the migration. Such a situation will be referred below simply to as electrophoresis. This is certainly not always the case, as the EDL symmetry may be distorted during the migration. This effect that is termed relaxation gives rise to an additional (retarding) field, which is directed oppositely to $E$. Note that (as a second order effect) the relaxation retardation, if any, is smaller than the one, caused by the motion of the ions of the undisturbed cloud. Since Onsager it is accepted that to arrive at a complete description of migration of ions, it is necessary to include this relaxation effect [9]. Note, however, the correction for relaxation goes to zero for large colloidal particles and is never included into consideration [28].
III Results and discussion
@@
We return to the issue of importance of the relaxation later, by focusing first on electrophoresis solely. Recall that this (leading order) contribution to a conductivity is traditionally calculated assuming that the clouds remain undisturbed during the migration. The electrophoretic mobilities of ions can be defined using the known expression for an electro-osmotic mobility [21, 30] taken with the opposite sign
@@
where $Z_{\pm}$ are the electrokinetic or zeta potentials of cations and anions, and $\zeta_{\pm}=eZ_{\pm}/k_{B}T$. In such a definition the geometry factor is hidden inside $\zeta$. Note that by introducing the dimensionless velocity $v_{\pm}=\dfrac{4\pi\eta\ell_{B}}{eE}V_{\pm}$, one can rewrite (1) as $v_{\pm}=\zeta_{\pm},$ which points clearly that the electrophoretic speed of ions is set by their zeta potentials.
The expression for zeta potentials of spherical isolating particles, which obey hydrodynamic no-slip boundary condition, has been derived already by Henry [31], and for ions of radius $R_{\pm}$ this can be formulated as
[…]
In order to take the integrals in Eq. (8) it is necessary to introduce explicitly the dependence of potentials on $r$. The $\phi_{\pm}$-profiles (and, consequently, $\phi_{s\pm}$) can be computed from Eq. (6) leading then to a numerical solution for $\zeta_{\pm}$ of ions.
When the linear approximation of Eq. (6) can be applied the local and surface potentials are given, approximately, by
$\phi_{\pm}\simeq\pm\dfrac{\ell_{B}}{\left(1+\varrho_{\pm}\right)}\dfrac{\mathrm{e}^{\varrho_{\pm}-r/\lambda_{D}}}{r},\,\phi_{s\pm}\simeq\pm\dfrac{\ell_{B}}{R_{\pm}\left(1+\varrho_{\pm}\right)}.$ (9)
It is normally considered that this is justified provided $|\phi_{s\pm}|\leq 1$, but we have recently shown that if radii are small, the linearization of Eq. (6) becomes no longer specific just to low potentials [22]. Thus Eqs. (9) should be a sensible approximation for inorganic salt ions of $|\phi_{s\pm}|=O(1)$.
Substituting equation (9) into (8) and performing the integration yields
@@
where $\mathcal{F_{\pm}}$ are the Henry special functions for nonconducting spherical particles [31]
@@
Figure 3: Function ±$\mathcal{F_{\pm}}/(1+\rho_{\pm})$ vs $\rho_{\pm}$ with $\mathcal{F_{\pm}}$ calculated from Eqs. (11) (solid curve), (12) (dashed curve), and ±$\mathcal{F_{\pm}}=2/3$ (dash-dotted curve). Dotted curve shows the electrophoretic term derived in [14].
From Eqs. (10) and (11) if follows that the salt dependence of $\zeta_{\pm}$ is hidden only in ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$. This dependence is illustrated in Fig. 3, where we plot ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ as a function of $\varrho_{\pm}$. The calculations are first made using $\mathcal{F_{\pm}}$ given by Eq. (11). It can be seen that the function ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ in Eq. (10) decreases [and vanishes in the limit $\lambda_{D}\to 0$]. Thus on increasing $\varrho_{\pm}$ (reducing $\lambda_{D}$) the migration speed of ions reduces, which is self-evident since the electrophoresis takes its origin in the EDL.
For small $\varrho_{\pm}$ Eq. (11) can be expanded about $\varrho_{\pm}=0$ and reduces to
±$\mathcal{F_{\pm}}\simeq\dfrac{2}{3}\left[1+\left(\dfrac{\varrho_{\pm}}{4}\right)^{2}\right],$ (12)
which has been used to derive the approximate conductivity equation [22]. To examine its range of applicability the curve for ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ calculated with $\mathcal{F_{\pm}}$ given by Eq. (12) is included in Fig. 3. The accuracy is quite good for $\varrho_{\pm}\leq 1$, but at larger salt concentrations there is some discrepancy in the direction of larger ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ than predicted by using Eq. (11). Moreover, the function exhibits its minimum at $\varrho_{\pm}\simeq 3$ and then augments, although can only decrease. Note, however, that the electrophoretic term derived by Avni et al. [14],
±$\frac{\mathcal{F_{\pm}}}{1+\varrho_{\pm}}\simeq\frac{2}{3}\left(1-\varrho_{\pm}e^{-a/\lambda_{D}}\right),$ (13)
where $a$ is the is the cut-off length defined as the sum of the cation and anion crystallographic radii, fails at much smaller $\varrho_{\pm}$. The calculation from Eq. (13) is included in Fig. 3. It can be seen that it shows significant deviations [from ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ obtained using (11)] already at $\varrho_{\pm}\geq 2\times 10^{-1}$ and turns to 2/32/32 / 3 when $\varrho_{\pm}\geq 2$. This implies that the zeta potentials become the same as at an infinite dilution, which is counterintuitive.
Also included in Fig. 3 is the curve calculated using ±$\mathcal{F_{\pm}}=2/3$. It seems clear that this provides ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ that practically coincides (when $\varrho_{\pm}\leq 1$) or is very close [at $\varrho_{\pm}=O(1)$] to obtained with Eq. (11). This implies that the decrease in $|\zeta_{\pm}|$ with salt is to the first order caused only by the reduction of the ion surface potential:
$\zeta_{\pm}\simeq\frac{2}{3}\phi_{s\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}(1+\varrho_{\pm})}.$ (14)
The upper bound for the zeta potential is attained when $\rho_{\pm}\to 0$ (an infinite dilution)
@@
Figure 4: Zeta potentials $\zeta_{\pm}$ vs. $c_{\infty}$ computed for K+, Li+, Cl-, I- (solid curves from top to bottom). Dashed lines are obtained using Eq. (15). Symbols show calculations from Eq. (14).
Figure 4 includes computed zeta potentials of several ions as a function of salt concentration. For these examples we take from Table 1 cations (K+, Li+) and anions (I-, Cl-) of the smallest and largest radius. As a side note, these numerical calculations [based on Eqs. (8) and (6)] fully coincide with those made from the first equality in (10) using $\phi_{s\pm}$ computed from the nonlinear Poisson-Boltzmann equation and $\mathcal{F_{\pm}}$ given by Eq. (11). The straight lines corresponding to $\zeta_{0\pm}$ given by (15) describe perfectly the distinct plateau regions at low concentrations. On increasing concentration, however, the absolute values of $\zeta_{\pm}$ decrease. Also included are calculations from Eq. (14). It can be seen that their fit to numerical data is extremely good in the whole concentration range.
Figure 5: $K/K_{0}$ as a function of $c_{\infty}$ for KBr, NaCl and LiI aqueous solution calculated from Eqs. (17), (18) and (19) (solid, dash-dotted and dashed curves). Open and filled circles indicate experimental data from [33] and [34], open and filled squares show data from [35] and [36].
@@
where $M_{0}$ and $K_{0}$ are given by Eqs. (4).
Figure 5 shows $K/K_{0}$ calculated from Eq. (17) as a function of $c_{\infty}$. The calculations are compared with experimental data for three standard salts [33, 34, 35, 36]. As seen from Table 1, for these examples $N$ varies from 1.1 for KBr, but can be larger than 2 (e.g. for LiI). An overall conclusion from this linear scale plot is that the theoretical curve for $K/K_{0}$ is well consistent with experiment. Only for KBr Eq.(17) predicts a slightly lower conductivity at large concentrations, but still fits data quite well. This is a startling result since our mean-field theory omits nonidealities that are normally considered to be important at high salt [and at this point does not include the relaxation effect]. Thus, our results suggest that, if any, they constitute a second-order correction, which could make only a small improvement to the data description.
When $N=1$, we get $\varrho_{h}=R_{h}/\lambda_{D}=2\varrho_{+}/(1+N)=\varrho_{+}$, and Eq. (17) reduces to
@@
Eq. (18) predicts a monotonic decrease in relative (electrophoretic) conductivity with $\varrho_{h}$ as it should be, and $K/K_{0}\to 0$ as $\varrho_{h}\to\infty$. For $\varrho_{h}\ll 1$ it reduces to the famous Onsager formula for the electrophoretic effect